OpenStudy (anonymous):

Calc Q: given a fxn f(x)=x^3+ax^2+bx+k, where a,b,and k are constants. The fxn has a local min at x=-1 and an inflection point at x=-2. a. find the values of a and b b. if the integral of f(x)dx=32 (from 0 to 1), what is the value of k? just need a general idea of process.

myininaya (myininaya):

find f' \[f'(x)=3x^2+2ax+b\] Since x=-1 is a min then f'(-1)=0 so we have \[f'(-1)=3+2a(-1)+b=0=> -2a+b=-3\] find f'' \[f''(x)=6x+2a\] Since x=-2 we have an inflection point then f''(-2)=0 so we have \[f''(-2)=6(-2)+2a=0 => 2a=12=>a=6\] if a=6 then -2(6)+b=-3 => -12+b=-3 => b=-3+12=9

myininaya (myininaya):

\[ \int\limits_{0}^{1}(x^3+6x^2+9x+k) dx=32\]

OpenStudy (anonymous):

Thank you!!